OpenStudy (anonymous):

PLEASE HELP Solve: | m + 8| = 16 A. {8} B. {-24, 8} C. {8, 16} D. {-24}

OpenStudy (anonymous):

Well 8+8 = 16 is the obvious. But m + 8 = -16 is also an solution since |x| = |-x| Can you find any m that satisfies the equation m+8 = -16